A body of mass m accelerates uniformly from rest to a speed v in time t , if the body is performing one-D motion then the work done on the body till any time T between 0 and t is:

Grade:12

1 Answers

Vikas TU
6 years ago
Dear Student,
  1. From the equation of motion,
    v=u+at (u=0,v=v,a=acceleration,t=t)
    =>v=at
    =>a=v/t
    Force=ma
    =mv/t
    Distance travelled in time T = S = ut+at2/2 (from the equations of motion)
    =>S=vT2/2t (since t=T, u=0, a=v/t)
    Work done=ForcexDistance travelled in that time T
    =mv/t x vT2/2t
    =mv2T2/2t2.
    Cheers!!
    Regards,
    Vikas (B. Tech. 4th year
    Thapar University)